# Notions of entropy for unitary representations

### v2

@@ 6.2 Upper bound and completed proof of Theorem A
Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be bounded, let $a:\Gamma\to\mathrm{M}_{k}$ be finitely supported, and let $(F_{n})_{n\geq 1}$ be a right Følner sequence. There are subsets $E_{n}$ of $F_{n}$ such that $|F_{n}\setminus E_{n}|=o(|F_{n}|)$ and
@@
vanishes in all columns indexed by $E_{n}$. On the other hand, all of its entries can be bounded by (5.2). The result is that
@@
Finally, every diagonal block of $(\varphi\ast a)[F_{n}]$ is simply equal to $(\varphi\ast a)(e)$. ∎
@@
The earliest result I know that includes Proposition 6.6 for any amenable group is [Den06, Theorem 3.2]; see also the theorem attributed to Lück and Schick that precedes it in that paper. The proof starts with estimates similar to Lemma 6.4, but in which all functions on $\Gamma$ are finitely supported. Using these, one shows by induction on $d$ that
@@
By Corollary 6.5 and the calculation (5.5), that normalized trace converges to
@@ 6.3 Further remarks
  • The proof of Theorem S presented in [Sim05a, Section 2.3], which is modeled on Verblunsky’s from [Ver35b, Ver35a], uses a classical cousin of the variational principle from Proposition 3.9. However, in [Sim05a, Section 2.3] that variational principle is used only to establish weak upper semicontinuity of the right-hand side of (1.1) as a function of $\mu$: see [Sim05a, Theorem 2.3.4]. This is a preparation for that proof of Theorem S, rather than an application inside the proof itself. In other respects, that proof of Theorem S seems essentially disjoint from our proof of Theorem A.
@@
On the other hand, as discussed in the Introduction, most traditional proofs of Theorem S make rather explicit use of a notion of the ‘past’ of a vector under a unitary operator. In the first place, this refers to the subspace $M$ in (1.4). In the setting of Theorem A we must do without this structure. (A meaning of ‘past’ re-appears when we turn to Theorem B in the next section.)
@@
In this respect, we take inspiration from research in ergodic theory that studies entropy without conditioning on the past as in (1.5). This program is discussed more fully in [GTW00]. Regarding Fuglede–Kadison determinants and generalizations of Theorem S itself, some other recent precedents for our work in ergodic theory also have this flavour.
@@
For a polynomial $f$ in $d$ variables, a log-integral much like the right-hand side of (1.1) defines its ‘Mahler measure’. If $f$ has integer coefficients, then it can be used to construct an action of $\mathbf{Z}^{d}$ by automorphisms of a compact Abelian group. This action necessarily preserves the Haar measure of that compact group, and the Kolmogorov–Sinai entropy of this measure-preserving action turns out to equal the logarithm of the Mahler measure of $f$: see [Sch95, Chapter V] and the references given there.
@@
Starting from the observation that a Mahler measure is a Fuglede–Kadison determinant, Deninger generalized this construction to allow $f\in\mathrm{M}_{k}(\mathbf{Z}[\Gamma])$ for any positive integer $k$ and countable group $\Gamma$. He conjectured that the Kolmogorov–Sinai entropy of the resulting system should be the Fuglede–Kadison determinant of $f$ whenever $\Gamma$ is amenable and $f$ is non-singular as an operator on $\ell^{2}(\Gamma)^{\oplus k}$. He proved this under various extra hypotheses: see [Den06, Den09, Den12] and also the joint work [DS07].
@@
Deninger proved [Den06, Theorem 3.2] for application in this context. This theorem contains our Proposition 6.6, and it actually allows any positive invertible element of $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$, although Deninger required additional hypotheses before reaching his main result in ergodic theory. As Deninger describes, several aspects of this proof already appear in earlier works on $L^{2}$-invariants in algebraic topology, such as [Lüc94, DM98, Sch01].
@@
Li made further progress in [Li12], which proved Deninger’s conjectured entropy formula whenever $\Gamma$ is amenable and $f$ is invertible in $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$. Among Li’s technical ingredients, [Li12, Corollary 7.2] strengthens Proposition 6.6 by allowing certain additional perturbations to each of the matrices $\varphi[F_{n}]$.
@@
Finally, Li and Thom proved Deninger’s full conjecture in [LT14], and generalized it further still to systems constructed by Pontrjagin duality from any $\mathbf{Z}[\Gamma]$-module of type FL. This work needs another strengthening of Proposition 6.6 as an ingredient: [LT14, Theorem 1.4], which allows positive elements of $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$ that are not necessarily invertible. This is equivalent to the case of Theorem A when $\varphi$ is $\lambda$-normal and the operator $T$ from Proposition 4.7 is bounded.
@@
Apart from these points of contact around [Den06, Theorem 3.2], the other details of our proof of Theorem A are largely disjoint from those previous works. The most obvious difference is our use of the variational principle from Proposition 4.10. In addition, not all aspects of Szegő’s theorem make an appearance in those previous papers: for example, they all assume that $f$ already lies in $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$, so there is no singular part to take care of.
@@
It is natural to ask about generalizations of Theorem A to non-amenable groups or other C*-algebras. Various directions suggest themselves, but none of them seems entirely straightforward. We quickly discuss three of these here. For simplicity, we now restrict our attention to scalar-valued positive definite functions and positive functionals.
@@
If this implication does hold, but $\Gamma$ is not amenable, then it might be worth looking more closely at the ‘gap’ in (6.8), perhaps for special elements of $\mathbf{C}[\Gamma]$ such as the Laplacians corresponding to finite symmetric subsets of $\Gamma$.
@@
For particular choice of $a$ as in Problem 6.7, how does the gap in (6.8) relate to other measures of non-amenability such as the isoperimetric profile of finite subsets of $\Gamma$?
@@
For our second direction, let us consider the possibility of replacing finite subsets of $\Gamma$ in Theorem A with finite quotients. This moves us to the class of residually finite groups. By allowing quotients with small defects, one could extend this further to sofic groups [Gro99, Wei00], but we leave these aside here.
If $\Gamma$ is residually finite, then it has a sequence of permutation representations $\sigma_{n}:\Gamma\to\mathrm{Sym}(V_{n})$ on finite sets such that the kernels $\ker\sigma_{n}$ have trivial intersection. In fact we may assume further that this sequence is asymptotically free, meaning that
@@
$(\det\pi_{n}(a))^{1/d_{n}}\to\Delta a\qquad\hbox{as}\qquad n\to\infty.$ (6.10)
The proof is closely analogous to that of Proposition 6.6. See [DS07, Theorem 6.1] (which makes the slightly stronger assumption that $a$ is invertible in the Banach algebra $\ell^{1}(\Gamma)$) and [KL11, Lemma 7.2 and Theorem 7.3].
@@
    Suppose that $a$ is non-negative but not necessarily invertible. The proof of (6.10) depends on approximating $\log$ by polynomials, and such approximations break down near the origin. In addition, our proof of Proposition 6.1 considers alternating products in the fixed Hilbert space of the regular representation, and has no obvious modification for the sequence of spaces $\mathbf{C}^{\oplus V_{n}}$. As a result, we now obtain only the inequality
@@
    It turns out that this inequality really can be strict [Lüc02, Example 13.69]. This happens if $a$ is non-singular and satisfies $\Delta a>-\infty$, but the finitary matrices $\pi_{n}(a)$ have a few extremely small eigenvalues that drag their determinants down far below $\Delta a$.
    Despite this possibility, the inequality above is still enough for some valuable applications. An example is Elek and Szabó’s proof of Lück’s determinant conjecture for sofic groups in [ES05]; see also [BS16].
@@
Hayes discusses the first of these difficulties further near the end of [Hay16, Introduction]. His work concerns an analog of Deninger’s entropy calculation for certain dynamical systems of algebraic origin over sofic groups. For Hayes, the change of focus to sofic entropy means that he can sidestep this first difficulty, and the second does not arise because sofic entropy is not defined directly as a limit of normalized Shannon entropy values. We meet a similar situation in our study of almost periodic entropy and Theorem C below. See also the discussion in Subsection 9.7.
@@
Sequences $(V_{n})_{n\geq 1}$ satisfying the two properties listed above have long-standing importance in the study of hyperfiniteness [Dix81, Chapter III.7]. This property is an abstraction of amenability for von Neumann algebras. Indeed, among factors of type II1 the existence of such a sequence identifies uniquely the hyperfinite one: see [Con76, Theorem 5.1] and also the related discussion in [Arv77, Section 1].
Beyond this simple result for von Neumann algebras, one could try to push Theorem A towards a class of abstract C*-algebras that satisfy some generalization of amenability. Two possibilities are nuclearity [Lan82] or quasidiagonality [Bro04], although both of these properties capture the amenability of a group $\Gamma$ through the reduced C*-algebra $\lambda(C^{\ast}\Gamma)$ rather than $C^{\ast}\Gamma$. However, I do not know precisely what statement one should try to prove.
@@ 7 Subdiagonal subalgebras and random orders
This section proves Theorem B. Let $\Gamma$ be a countable group, $\lambda$ be its left regular representation, and $\xi$ be the usual cyclic tracial unit vector for $\lambda$.
In large part, Theorem B is a special case of the Arveson–Labuschagne generalized Szegő theorem for subdiagonal subalgebras. We recall the necessary definitions and this theorem in Subsection 7.5.
However, some work is required to make contact with that theorem. In the first place, Arveson’s subdiagonal subalgebras are utraweakly closed subalgebras of von Neumann algebras, not subalgebras of abstract C*-algebras. This setting provides some of the auxiliary objects required for their definition, starting with a compatible conditional expectation operator. So we need to put our objects into this form.
If $\varphi$ is $\lambda$-normal, then we do this by enlarging $\lambda(\Gamma)^{\prime\prime}$ to a von Neumann crossed product algebra that ‘knows’ the law of a translation-invariant random total order. More generally, we must first prove directly that $\varphi_{\mathrm{sing}}$ makes no contribution to the left-hand side of (1.7), thereby reducing our work to the $\lambda$-normal case.
@@
Similar uses of crossed product von Neumann algebras have appeared previously in the study of random processes on Cayley graphs or unimodular random networks. See, for example, [Gab05, AL07] and the references given there. However, as far as I know, the connection to subdiagonal subalgebras is new.
The study of subdiagonal subalgebras belongs firmly within the abstract theory of von Neumann algebras. As a result, this section cites some more advanced results from the literature of that field than the rest of the paper.
@@ 7.1 Background on subdiagonal subalgebras
The inclusion of the Hardy space $H^{\infty}(\mathbf{T})$ in $L^{\infty}(\mathbf{T})$ contains those functions whose Fourier transform is one-sided. These can be interpreted as the functions that ‘depend only on the past’ in Fourier space. In [Arv67], Arveson defined a subdiagonal subalgebra of a finite von Neumann algebra $\mathfrak{M}$ by axioms abstracted from that example. It can be seen as providing $\mathfrak{M}$ with an abstract notion of ‘past’.
To formulate Arveson’s definition, let $\Phi$ be a faithful normal conditional expectation from $\mathfrak{M}$ onto a von Neumann subalgebra of $\mathfrak{M}$, in the sense studied by Dixmier [Dix53], Umegaki [Ume54, Ume56, Ume59] and Tomiyama [Tom57, Tom58, Tom59]; see also [Arv67, Definition 2.1.3 and Chapter 6].
The following combines [Arv67, Definitions 2.1.1 and 2.2.2] (and uses slightly different notation). We also restrict our attention to unital subalgerbas for simplicity, although [Arv67] does not.
@@
Note that $\mathfrak{H}$ need not be closed, and in non-trivial cases it is not invariant under taking adjoints. By [Arv67, Theorem 2.2.1], any subdiagonal subalgebra with respect to $\Phi$ is contained in exactly one maximal subdiagonal subalgebra.
@@
A maximal subdiagonal subalgebra is necessarily ultraweakly closed, but the reverse implication is not known to hold in general: see [Arv67, Remark 2.2.3]. However, it does hold under hypotheses that cover our present needs.
@@
If $\mathfrak{H}_{0}$ is ultraweakly closed, then it is equal to $\mathfrak{H}$. In general, $\mathfrak{H}_{0}$ (resp. $\mathfrak{J}_{0}$) is ultraweakly dense in $\mathfrak{H}$ (resp. $\mathfrak{J}$). ∎
The first part of Theorem 7.2 is a theorem of Exel [Exe88, Theorem 7], subsequently generalized in [Ji00, Xu05]. The conclusion about $\mathfrak{H}_{0}$ and $\mathfrak{H}$ follows from this because it implies that the ultraweak closure of $\mathfrak{H}_{0}$ is already maximal. Finally, the conclusion about $\mathfrak{J}_{0}$ and $\mathfrak{J}$ follows because
@@
In the form of (7.1), this theorem was proposed in [Arv67], and proved there in special cases. The general case of (7.1) was proved much more recently in [Lab05] in the equivalent form of an abstract ‘Jensen formula’. Note that those references define the Fuglede–Kadison determinant of a functional via the variational principle, but this is equivalent to our formulation because of Proposition 3.9.
@@
A modern introduction to conditional expectations between C*-algebras can be found in [BO08, Section 1.5]. These maps find several applications later in that book.

### v3

@@ 6.2 Upper bound and completed proof of Theorem A
Let $\varphi:\Gamma\to\mathrm{M}_{k}$, let $a:\Gamma\to\mathrm{M}_{k}$ be finitely supported, and let $(F_{n})_{n\geq 1}$ be a right Følner sequence. There are subsets $E_{n}$ of $F_{n}$ such that $|F_{n}\setminus E_{n}|=o(|F_{n}|)$ and
@@
vanishes in all columns indexed by $E_{n}$. On the other hand, all of its entries can be bounded using (5.2). The result is that
@@
Finally, every diagonal block of $(\varphi\ast a)[F_{n}]$ is simply equal to $(\varphi\ast a)(e)$, so
$\overline{\mathrm{tr}}_{k|F_{n}|}\big{(}(\varphi\ast a)[F_{n}])=\overline{\mathrm{tr}}_{k}((\varphi\ast a)(e)).$
@@
The earliest reference I know that includes Proposition 6.6 for general amenable groups is [43, Theorem 3.2]; see also the theorem attributed to Lück and Schick that precedes it in that paper. The proof starts with estimates similar to Lemma 6.4, but in which all functions on $\Gamma$ are finitely supported. Using these, one shows by induction on $d$ that
@@
By Corollary 6.5 and the calculation (5.5), the last normalized trace converges to
@@ 6.3 Further remarks
  • The proof of Theorem S presented in [136, Section 2.3], which is modeled on Verblunsky’s from [152, 151], uses a classical cousin of the variational principle from Proposition 3.9. However, in [136, Section 2.3] that variational principle is used only to establish weak upper semicontinuity of the right-hand side of (1.1) as a function of $\mu$: see [136, Theorem 2.3.4]. This is a preparation for that proof of Theorem S, but not really an application inside the proof itself. In other respects, that proof of Theorem S seems essentially disjoint from our proof of Theorem A.
@@
On the other hand, as discussed in the Introduction, most traditional proofs of Theorem S make rather explicit use of a notion of the ‘past’ of a vector under a unitary operator. In the first place, this refers to the subspace $M$ in (1.4). In the setting of Theorem A we must do without this structure. A meaning of ‘past’ reappears when we turn to Theorem B in the next section.
@@
In this respect, we take inspiration from research in ergodic theory that studies entropy without conditioning on the past as in (1.5). This program is discussed more fully in [71]. Regarding Fuglede–Kadison determinants and generalizations of Theorem S itself, some other recent precedents for our work in ergodic theory also have this flavour.
@@
For a polynomial $f$ in $d$ variables, a log-integral much like the right-hand side of (1.1) defines its ‘Mahler measure’. If $f$ has integer coefficients, then it can be used to construct an action of $\mathbf{Z}^{d}$ by automorphisms of a compact Abelian group. This action necessarily preserves the Haar measure of that compact group, and the Kolmogorov–Sinai entropy of this measure-preserving action turns out to equal the logarithm of the Mahler measure of $f$: see [130, Chapter V] and the references given there.
@@
Starting from the observation that a Mahler measure is a Fuglede–Kadison determinant, Deninger generalized this construction to allow $f\in\mathrm{M}_{k}(\mathbf{Z}[\Gamma])$ for any positive integer $k$ and countable group $\Gamma$. He conjectured that the Kolmogorov–Sinai entropy of the resulting system $\boldsymbol{X}_{f}$ should be the Fuglede–Kadison determinant of $f$ whenever $\Gamma$ is amenable and $f$ is non-singular as an operator on $\ell^{2}(\Gamma)^{\oplus k}$. He proved this under various extra hypotheses: see [43, 44, 45] and also the joint work [46].
@@
Deninger proved [43, Theorem 3.2] in order to apply it in this context. This theorem contains our Proposition 6.6, and it actually allows any positive invertible element of $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$, although Deninger required additional hypotheses before reaching his main result in ergodic theory. As Deninger describes, several aspects of this proof already appear in earlier works on $L^{2}$-invariants in algebraic topology such as [106, 50, 129].
@@
Li made further progress in [103], where he proved Deninger’s conjectured entropy formula whenever $\Gamma$ is amenable and $f$ is invertible in $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$. Among Li’s technical ingredients, [103, Corollary 7.2] strengthens Proposition 6.6 by allowing certain additional perturbations to each of the matrices $\varphi[F_{n}]$.
@@
Finally, Li and Thom proved Deninger’s full conjecture in [104], and generalized it further still to systems constructed by Pontrjagin duality from any $\mathbf{Z}[\Gamma]$-module of type FL. This work needed another strengthening of Proposition 6.6 as an ingredient: see [104, Theorem 1.4], which allows positive elements of $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$ that are not necessarily invertible. This is equivalent to the case of Theorem A when $\varphi$ is $\lambda$-normal and the operator $T$ from Proposition 4.10 is bounded.
@@
Apart from these points of contact around [43, Theorem 3.2], the other details of our proof of Theorem A are largely disjoint from those previous works. The most obvious difference is our use of the variational principle from Proposition 4.13. In addition, not all aspects of Szegő’s theorem make an appearance in those previous papers: for example, they all assume that $f$ already lies in $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$, so there is no singular part to take care of.
@@
It is natural to ask about generalizations of Theorem A to non-amenable groups or other C*-algebras. Various directions suggest themselves, but none of them seems entirely straightforward. We quickly discuss four of these here. For simplicity, we now restrict our attention to scalar-valued positive definite functions and positive functionals.
@@
If the answer here is positive and $\Gamma$ is not amenable, then it might be worth looking more closely at the ‘gap’ in (6.8), perhaps for special elements of $\mathbf{C}[\Gamma]$ such as the Laplacians corresponding to finite symmetric subsets of $\Gamma$.
@@
For particular choices of $a$ as in Problem 6.7, how does the gap in (6.8) relate to other measures of non-amenability such as the isoperimetric profile of finite subsets of $\Gamma$?
@@
For our second direction, let us consider the possibility of replacing finite subsets of $\Gamma$ in Theorem A with finite quotients. This moves us to the class of residually finite groups. By allowing quotients with small defects, one could extend this further to sofic groups [76, 159], but we leave these aside here.
If $\Gamma$ is residually finite, then it has a sequence of permutation representations $\sigma_{n}:\Gamma\to\mathrm{Sym}(V_{n})$ on finite sets such that the kernels $\ker\sigma_{n}$ have trivial intersection. In fact, by replacing $\sigma_{n}$ with the diagonal action of many copies of $\sigma_{n}$ if necessary, we may assume further that this sequence is asymptotically free, meaning that
@@
$(\det\pi_{n}(a))^{1/|V_{n}|}\to\Delta a\qquad\hbox{as}\qquad n\to\infty.$ (6.10)
The proof is closely analogous to that of Proposition 6.6. See [46, Theorem 6.1] (which makes the slightly stronger assumption that $a$ is invertible in the Banach algebra $\ell^{1}(\Gamma)$) and [95, Lemma 7.2 and Theorem 7.3].
@@
    Suppose that $a$ is non-negative but not necessarily invertible. The proof of (6.10) depends on approximating $\log$ by polynomials, and such approximations break down near the origin. On the other hand, our proof of Proposition 6.1 considers alternating products in the fixed Hilbert space of the regular representation, and has no obvious modification for the sequence of spaces $\mathbf{C}^{\oplus V_{n}}$. Absent either of these arguments, we now obtain only the inequality
@@
    It turns out that this inequality really can be strict [107, Example 13.69]. This happens if $a$ is non-singular and satisfies $\Delta a>-\infty$, but the finitary matrices $\pi_{n}(a)$ have a few extremely small eigenvalues that drag their determinants down far below $\Delta a$.
    Despite this possibility, the inequality above is still enough for some valuable applications. An example is Elek and Szabó’s proof of Lück’s determinant conjecture for sofic groups in [53]; see also [13].
@@
Hayes discusses the first of these difficulties further near the end of [80, Introduction]. That paper concerns an analog of Deninger’s entropy calculation for certain dynamical systems of algebraic origin over sofic groups. For Hayes, the change of focus to sofic entropy meant that he could sidestep the first difficulty above, and the second did not arise because sofic entropy is not defined directly as a limit of normalized Shannon entropy values. We meet a similar situation in our study of almost periodic entropy and Theorem C below. See also the discussion in Subsection 9.7.
@@
Sequences $(V_{n})_{n\geq 1}$ satisfying the two properties above have long-standing importance in the study of hyperfiniteness [49, Chapter III.7]. This property is an abstraction of amenability for von Neumann algebras. Indeed, among factors of type II1, the existence of such a sequence identifies uniquely the hyperfinite one: see [36, Theorem 5.1] and also the related discussion in [8, Section 1].
Beyond this simple result for von Neumann algebras, one could try to push Theorem A towards a class of abstract C*-algebras that satisfy some generalization of amenability. Two possibilities are nuclearity [101] or quasidiagonality [29], although both of these properties capture the amenability of a group $\Gamma$ through the reduced C*-algebra $\lambda(C^{\ast}\Gamma)$ rather than $C^{\ast}\Gamma$. However, I do not know precisely what statement one should try to prove, and I suspect that versions of difficulties (1) and (2) above could re-emerge.
@@
Finally, one could look for a version of Theorem A in a setting that generalizes amenable groups from a more combinatorial point of view. For example, amenability is defined for unimodular random networks in [3, Section 8]. With the right choice of C*-algebra to consider in place of the group C*-algebra, I expect a generalization of Theorem A in this direction could be fairly straightforward.
@@ 7 Subdiagonal subalgebras and random orders
This section proves Theorem B. Let $\Gamma$ be a countable group, $\lambda$ be its left regular representation, and $\xi$ be the usual cyclic tracial unit vector for $\lambda$. Let $\varphi$ be an $\mathrm{M}_{k}$-valued positive definite function on $\Gamma$.
In large part, Theorem B is a special case of the Arveson–Labuschagne generalized Szegő theorem for subdiagonal subalgebras. We recall this theorem in Subsection 7.1.
However, some work is required to make contact with that theorem. In the first place, Arveson’s subdiagonal subalgebras are utraweakly closed subalgebras of finite von Neumann algebras, not subalgebras of abstract C*-algebras, so we need to put our objects into this form.
If $\varphi$ is $\lambda$-normal, then we do this by enlarging $\lambda(\Gamma)^{\prime\prime}$ to a von Neumann crossed product algebra that ‘knows’ the law of some translation-invariant random total order. In general, we must first prove directly that $\varphi_{\mathrm{sing}}$ makes no contribution to the left-hand side of (1.7), thereby reducing our work to the $\lambda$-normal case.
@@
Similar uses of crossed product von Neumann algebras have appeared previously in the study of random processes on Cayley graphs or unimodular random networks. See, for example, [66, 3] and the references given there. However, as far as I know, the connection to subdiagonal subalgebras is new.
The study of subdiagonal subalgebras belongs firmly within the abstract theory of von Neumann algebras. As a result, this section cites some more advanced results from the literature of that field than the rest of this paper.
@@ 7.1 Background on subdiagonal subalgebras
Within $L^{\infty}(\mathbf{T})$, the Hardy space $H^{\infty}(\mathbf{T})$ contains those functions whose Fourier transform is one-sided. These can be interpreted as the functions that ‘depend only on the past’ in Fourier space. In [9], Arveson defined a subdiagonal subalgebra of a finite von Neumann algebra $\mathfrak{M}$ by axioms abstracted from this example. It can be seen as providing $\mathfrak{M}$ with an abstract notion of ‘past’.
To formulate Arveson’s definition, let $\Phi$ be a faithful normal conditional expectation from $\mathfrak{M}$ onto a von Neumann subalgebra of $\mathfrak{M}$, in the sense studied by Dixmier [47], Umegaki [147, 148, 149] and Tomiyama [144, 145, 146]; see also [9, Definition 2.1.3 and Chapter 6].
The following combines [9, Definitions 2.1.1 and 2.2.2] (and uses slightly different notation). We also restrict our attention to unital subalgerbas for simplicity, although [9] does not.
@@
Note that $\mathfrak{H}$ need not be closed, and in non-trivial cases it is not invariant under taking adjoints. By [9, Theorem 2.2.1], any subdiagonal subalgebra with respect to $\Phi$ is contained in exactly one maximal subdiagonal subalgebra.
@@
A maximal subdiagonal subalgebra is necessarily ultraweakly closed, but the reverse implication is not known to hold in general: see [9, Remark 2.2.3]. However, it does hold under hypotheses that cover our present needs.
@@
If $\mathfrak{H}_{0}$ is ultraweakly closed, then it is equal to $\mathfrak{H}$. In general, $\mathfrak{H}_{0}$ (respectively, $\mathfrak{J}_{0}$) is ultraweakly dense in $\mathfrak{H}$ (respectively, $\mathfrak{J}$). ∎
The first conclusion of Theorem 7.2 is a theorem of Exel [55, Theorem 7], subsequently generalized in [89, 160]. The conclusion about $\mathfrak{H}_{0}$ and $\mathfrak{H}$ follows from this because it implies that the ultraweak closure of $\mathfrak{H}_{0}$ is already maximal. Finally, the conclusion about $\mathfrak{J}_{0}$ and $\mathfrak{J}$ follows because
@@
In the form of (7.1), this theorem was proposed in [9], and proved there in special cases. The general case of (7.1) was proved much more recently in [100] in the equivalent form of an abstract ‘Jensen formula’. Note that those references define the Fuglede–Kadison determinant of a functional via the variational principle, but this is equivalent to our formulation because of Proposition 3.9.
@@
A modern introduction to conditional expectations between C*-algebras can be found in [30, Section 1.5]. These maps find several applications later in that book.